# Exact diagonalization study of energy level statistics in harmonically confined interacting bosons

### v1

@@ Abstract
[…]
@@ I Introduction
The energy-level spectra of a system exhibits unique features intrinsic to the system. Random Matrix Theory (RMT) as a framework for analyzing the statistical properties of energy spectra in heavy nuclei was introduced by E. P. Wigner [1, 2, 3, 4]. It was further developed by Mehta and Dyson [5, 6, 7] and has since become an important tool in the statistical analysis of energy levels in complex quantum many-body systems. In order to compare various models with the predictions of RMT, it is essential to eliminate from the energy-level spectra, the system-specific properties. This procedure, commonly referred to in the literature as unfolding [8], normalizes the mean level density to unity. However, the unfolding process may introduce certain inaccuracies due to the extraction of system-specific features. Energy level statistics measures such as the nearest-neighbor spacing distribution (NNSD) $P(s)$ [5, 9] and the distribution of the ratio of consecutive level spacings $P(r)$ [10] are used to probe short-range correlations, while the Dyson-Mehta $\Delta_{3}$ statistic [5] and the level number variance $\Sigma^{2}(L)$ [11] provide insight into long-range correlations. The NNSD $P(s)$, the Dyson-Mehta $\Delta_{3}$ statistic and the level number variance $\Sigma^{2}(L)$ involve the unfolding of the energy-level spectra while in the distribution of the ratio of consecutive level spacings $P(r)$, the unfolding is not required. Consequently, the NNSD $P(s)$, the Dyson-Mehta $\Delta_{3}$ statistic and the level number variance $\Sigma^{2}(L)$ may be prone to errors. For this reason, the distribution of the ratio of consecutive level spacings $P(r)$ is more useful tool than the NNSD $P(s)$ for the short-range correlations.
@@
The Bohigas-Giannoni-Schmit (BGS) conjecture [12] states that the nearest-neighbor spacing distribution of the energy-level spectra of quantum systems whose classical counterpart is chaotic follows one of the random matrix ensembles: the Gaussian Orthogonal Ensemble (GOE) or the Gaussian Unitary Ensemble (GUE) or the Gaussian Symplectic Ensemble (GSE) as determined by the time-reversal symmetry of the Hamiltonian [6, 7]. On the other hand, the Berry-Tabor conjecture demonstrates that the quantum systems whose classical analog is integrable, the nearest-neighbor spacing distribution of energy spectra follows the Poisson statistics [13]. Both conjectures have received substantial theoretical and numerical support and serve as a foundation for characterizing quantum chaos. RMT has found applications across diverse fields, including the stock market [14, 15], the atmospheric science [16] and the analysis of human electroencephalogram (EEG) data [17], etc.
@@
The statistical properties of energy levels in interacting trapped bosons have been explored for varying numbers of energy levels [18], and for both the lowest and the highest-lying energy levels [19, 20, 21] for a given two-body interaction. These studies reveal the impact of the two-body interaction and the trap energy on the energy level statistics, as well as deviation from the BGS conjecture. These works employ the potential harmonic expansion method (PHEM), which is an approximate many-body technique for calculating the energy level spectra. To the best of our knowledge, the effect of variation in two-body interaction, the number of bosons and rotation on the energy level statistics in interacting trapped bosons has not been examined. The aim of our present work is to fill this gap. This study holds particular significance as the many-body energy levels obtained from the diagonalization of the Hamiltonian matrix are variationally exact. A recent experimental study on ultracold gas of erbium atoms has observed signatures of quantum chaos [22]. Spectral statistics has also been explored across a range of systems to investigate the emergence of quantum chaos in two-component interacting fermions in a double-well potential [23], van der Waals clusters [24], Bose-Bose mixtures [25] and Hubbard model [26, 27, 28, 29].
@@
Following the experimental realization of Bose-Einstein condensation (BEC) in dilute atomic vapors of 87Rb [30], 23Na [31] and 39K [32] with repulsive interaction, and 7Li with attractive interaction [33, 34], there has been a surge in research activity employing BEC systems as a testbed for quantum many-body theories. Notably, the creation of vortices in rotating BEC [35, 36] has drawn significant attention from the superfluidity [37, 38, 39] and superconductivity [40, 41] communities in condensed matter physics. The system of trapped interacting bosons with an externally impressed rotation, as described in Sec. II, has been the subject of extensive theoretical and experimental investigations in the context of BEC. Statistical fluctuations in trapped interacting bosons are governed by two-body interactions characterized by the s-wave scattering length $a_{s}$ [42, 43]. In these systems there exist two energy scales, namely the trap energy ($\hbar\omega_{\perp}$) and the interaction energy ($\propto Na_{s}$, $N$ being the number of bosons). The interplay between these energy scales gives rise to rich physics within the system. There has been studies of various properties of the ground state such as the condensate fraction, breathing modes, vortex formation and vortex melting [44, 45, 46, 47, 48, 49, 50]. Our purpose of the current work is to investigate the statistical properties of the energy levels including the ground state and the low-lying excited states of the trapped interacting bosons, with focus on the interplay between the trap energy and the interaction energy.
@@ II The Model Hamiltonian
We start with a system of interacting bosons, each of mass $M$, trapped in a quasi-2D plane by a harmonic potential $V(r)=\frac{1}{2}M(\omega_{\perp}^{2}r_{\perp}^{2}+\omega_{z}^{2}z^{2})$ with $x$-$y$ symmetry. The system is also subjected to an externally impressed rotation with angular velocity $\Omega\equiv\Omega\hat{z}$. The many-body Hamiltonian for the system in the co-rotating frame is given by [44]
$\hat{H}^{rot}=\hat{H}^{lab}-\Omega.\hat{L}^{lab}$
@@
$\begin{split}\hat{H}^{lab}=&\sum_{i=1}^{N}\left[\frac{1}{2M}\left(\frac{\hbar}{i}\nabla_{i}\right)^{2}+\frac{1}{2}M\omega_{\perp}^{2}(r_{\perp i}^{2}+\lambda_{z}^{2}z_{i}^{2})\right]\\ &+\frac{1}{2}\frac{4\pi\hbar^{2}a_{sc}}{M}\left(\frac{1}{\sqrt{2\pi}\sigma}\right)^{3}\\ &\times\sum_{i\neq j}\exp^{-(\frac{1}{2\sigma^{2}})\{(r_{\perp i}-r_{\perp j})^{2}+(z_{i}-z_{j})^{2}\}}\end{split}$ (1)
@@
$\hat{L_{z}}^{lab}=\sum_{i=1}^{N}\hat{l_{zi}}^{lab}=\frac{\hbar}{i}\sum_{i=1}^{N}(r_{i}\times\nabla_{i})_{z},$
$\hat{L_{z}}^{lab}$ is the total angular momentum of the system in the laboratory frame. Here $r_{\perp}=\sqrt{x^{2}+y^{2}}$ is the radial distance of the particle from the trap center; $\omega_{\perp}$ and $\omega_{z}$ are the confining harmonic frequencies in the $x$-$y$ plane and the axial direction, respectively. The first two terms of the many-body Hamiltonian in Eq. (1) represent the kinetic energy and the external confining potential, while the last term describes the two-body interaction, modeled by a potential Gaussian in particle-particle separation. We introduce the anisotropic parameter $\lambda_{z}\equiv\frac{\omega_{z}}{\omega_{\perp}}\gg$1 which signifies that the trap is highly oblate spheroidal and hence the confined system is effectively quasi-2D. We choose $\hbar\omega_{\perp}$ as the unit of energy and define $a_{\perp}$=$\sqrt{\frac{\hbar}{M\omega_{\perp}}}$ as the unit of length. In the Gaussian potential, $\sigma$ (scaled by $a_{\perp}$) is the interaction range of the potential. The two-body interaction strength after tracing out the $z$-coordinate becomes $g_{2}=\frac{4\pi\hbar^{2}a_{sc}}{M}\frac{1}{\sqrt{2\pi}}\sqrt{\frac{\lambda_{z}}{a_{\perp}^{2}[1+(\frac{\sigma}{a_{\perp}})^{2}\lambda_{z}]}}$ which in the dimensionless form turns out to be $g_{2}=\frac{4\pi a_{sc}}{a_{\perp}}\sqrt{\frac{\lambda_{z}}{2\pi}}$ in the limit $\frac{\sigma}{a_{\perp}}\ll 1$. Here $a_{sc}$ is the s-wave scattering length for the two-body interaction. Experimentally, one can vary the scattering length $a_{sc}$ employing the Feshbach resonance [42, 51]. In this study we assume $a_{sc}>0$ so that the effective interparticle potential is in the repulsive regime. When the interaction range approaches zero i.e $\sigma\to 0$, the Gaussian interaction reduces to the $\delta$-function potential $V(r_{\perp i},r_{\perp j})=g_{2}\delta(r_{\perp i}-r_{\perp j})$ [43].
The scheme for constructing many-body basis states for the model has been described in ref. [44].
@@ III.1 Nearest-neighbor spacing distribution
where $b$ is the Brody parameter and $a=[\Gamma(\frac{2+b}{1+b})]^{1+b}$ [$\Gamma$ is Euler’s gamma function]. The parameter $b$ is like an interpolation parameter which for $b=0$ corresponds to the Poisson distribution while $b=1$ corresponds to the GOE distribution.
@@ III.2 The Dyson-Mehta $\Delta_{3}(L)$ statistic
[…]
@@
$\langle\Delta_{3}(L)\rangle=\frac{L}{15}$
@@
$\langle\Delta_{3}(L)\rangle=\frac{1}{\pi^{2}}\ln(L)-0.00695.$

### v2

@@ Abstract
We present an exact diagonalization study of the spectral properties of bosons harmonically confined in a quasi-2D plane and interacting via repulsive Gaussian potential. We consider the lowest $100$ energy levels for systems of $N=12,16$ and $20$ bosons in two distinct regimes: (a) when the interaction energy is small compared to the trap energy (moderate interaction) and (b) when the interaction energy is comparable to the trap energy (strong interaction), for the non-rotating ($L_{z}=0$) as well as the rotating single-vortex state ($L_{z}=N$). For higher angular momenta, $L_{z}=2N$ and $L_{z}=3N$, only the strong interaction regime is considered. While the nearest-neighbor spacing distribution (NNSD) $P(s)$ and the ratios of consecutive level spacings distribution $P(r)$ are used to study the short-range correlations, the Dyson-Mehta $\Delta_{3}$ statistic and the level number variance $\Sigma^{2}(L)$ are used to examine the long-range correlations. In the moderate interaction regime, the non-rotating system exhibits Poisson distribution, a characteristic of the regular energy spectra. In the strong interaction regime, the non-rotating system exhibits chaotic behavior signified by GOE distribution. Furthermore, in the rotating case for the single-vortex state ($L_{z}=N$) in the moderate interaction regime, the system exhibits signatures of weak chaos with some degree of regularity in the energy-level spectra. However, in the strong interaction regime for the rotating case with $L_{z}=N$, $2N$ and $3N$, the system exhibits strong chaotic behavior. The rotation is found to contribute to enhancement of chaotic behavior in the system for both the moderate and the strong interaction regimes. Our results of NNSD analysis are supported by the analysis of the ratios of consecutive level spacings distribution $P(r)$, which does not involve unfolding.
@@ I Introduction
The energy-level spectra of a system exhibits unique features intrinsic to the system. Random Matrix Theory (RMT) as a framework for analyzing the statistical properties of energy spectra in heavy nuclei was introduced by E. P. Wigner [1, 2, 3, 4]. It was further developed by Mehta and Dyson [5, 6, 7] and has since become an important tool in the statistical analysis of energy levels in complex quantum many-body systems. In order to compare various models with the predictions of RMT, it is essential to eliminate from the energy-level spectra, the system-specific properties. This procedure, commonly referred to in the literature as unfolding [8], normalizes the mean level density to unity. While extracting system-specific features, the possible error arising out of the unfolding process can be minimized by fitting a low-degree polynomial or by performing linear fits over smaller intervals with moving average technique. Energy level statistics measures such as the nearest-neighbor spacing distribution (NNSD) $P(s)$ [5, 9, 10] and the distribution of the ratio of consecutive level spacings $P(r)$ [11] are used to probe short-range correlations, while the Dyson-Mehta $\Delta_{3}$ statistic [5] and the level number variance $\Sigma^{2}(L)$ [12] provide insight into long-range correlations. The NNSD $P(s)$, the Dyson-Mehta $\Delta_{3}$ statistic and the level number variance $\Sigma^{2}(L)$ involve the unfolding of the energy-level spectra while in the distribution of the ratio of consecutive level spacings $P(r)$, the unfolding is not required. Consequently, the NNSD $P(s)$, the Dyson-Mehta $\Delta_{3}$ statistic and the level number variance $\Sigma^{2}(L)$ may be prone to errors. For this reason, the distribution of the ratio of consecutive level spacings $P(r)$ is more useful tool than the NNSD $P(s)$ for the short-range correlations.
@@
The Bohigas-Giannoni-Schmit (BGS) conjecture [13] states that the nearest-neighbor spacing distribution of the energy-level spectra of quantum systems whose classical counterpart is chaotic follows one of the random matrix ensembles: the Gaussian Orthogonal Ensemble (GOE) or the Gaussian Unitary Ensemble (GUE) or the Gaussian Symplectic Ensemble (GSE) as determined by the time-reversal symmetry of the Hamiltonian [6, 7]. On the other hand, the Berry-Tabor conjecture demonstrates that the quantum systems whose classical analog is integrable, the nearest-neighbor spacing distribution of energy spectra follows the Poisson statistics [14]. Both conjectures have received substantial theoretical and numerical support and serve as a foundation for quantum chaos. RMT has found applications across diverse fields, including the stock market [15, 16], the atmospheric science [17] and the analysis of human electroencephalogram (EEG) data [18], etc.
@@
The statistical properties of energy levels in interacting trapped bosons have been investigated for varying numbers of energy levels [19], as well as for the lowest and the highest lying levels [20, 21, 22], for a given two-body interaction, using the approximate potential harmonic expansion method (PHEM). These works examine the effects of the two-body interaction and the trap energy on the energy level statistics, reporting deviation from the BGS conjecture. To the best of our knowledge, the effect of variation in the two-body interaction, the number of bosons and the rotation on the energy level statistics of trapped interacting bosons has not been explored. In the present study, we include rotation and extend the analysis to two distinct regimes: one in which the interaction energy is small compared to the trap energy (moderately interacting regime) and the other in which the interaction energy is comparable to the trap energy (strongly interacting regime). Our present study holds particular significance as the many-body energy levels obtained from the diagonalization of the Hamiltonian matrix are variationally exact. A recent experimental study on ultracold gas of erbium atoms has observed signatures of quantum chaos [23]. Spectral statistics has also been explored across a range of systems to investigate the emergence of quantum chaos, including two-component interacting fermions in a double-well potential [24], van der Waals clusters [25], Bose-Bose mixtures [26] and Hubbard model [27, 28, 29, 30].
@@
Following the experimental realization of Bose-Einstein condensation (BEC) in dilute atomic vapors of 87Rb [31], 23Na [32] and 39K [33] with repulsive interaction, and 7Li with attractive interaction [34, 35], there has been a surge in research activity employing BEC systems as a testbed for quantum many-body theories. Notably, the creation of vortices in rotating BEC [36, 37] has drawn significant attention from the superfluidity [38, 39, 40] and superconductivity [41, 42] communities in condensed matter physics. The system of trapped interacting bosons with an externally impressed rotation, as described in Sec. II, has been the subject of extensive theoretical and experimental investigations in the context of BEC. Statistical fluctuations in trapped interacting bosons are governed by two-body interactions characterized by the s-wave scattering length $a_{s}$ [43, 44]. In these systems there exist two energy scales, namely the trap energy ($\hbar\omega_{\perp}$) and the interaction energy ($\propto Na_{s}$, NN being the number of bosons). The interplay between these energy scales gives rise to rich physics within the system. There has been studies of various properties of the ground state such as the condensate fraction, breathing modes, vortex formation and vortex melting [45, 46, 47, 48, 49, 50, 51]. Our purpose of the current work is to investigate the statistical properties of the energy levels including the ground state and the low-lying excited states of the trapped interacting bosons, with focus on the interplay between the interaction energy and the trap energy.
@@ II The Model Hamiltonian
We start with a system of interacting bosons, each of mass $M$, trapped in a quasi-2D plane by a harmonic potential $V(r)=\frac{1}{2}M(\omega_{\perp}^{2}r_{\perp}^{2}+\omega_{z}^{2}z^{2})$ with x{x}-y{y} symmetry. The system is also subjected to an externally impressed rotation with angular velocity $\Omega\equiv\Omega\hat{z}$. The many-body Hamiltonian for the system in the co-rotating frame is given by [45]
$\hat{H}^{\mathrm{rot}}=\hat{H}^{\mathrm{lab}}-\Omega.\hat{L}^{\mathrm{lab}}$
@@
$\begin{split}\hat{H}^{\mathrm{lab}}=&\underbrace{\sum_{i=1}^{N}\left[\overbrace{\frac{1}{2M}\left(\frac{\hbar}{i}\nabla_{i}\right)^{2}+\frac{1}{2}M\omega_{\perp}^{2}(r_{\perp i}^{2}+\lambda_{z}^{2}z_{i}^{2})}^{\hat{h}(\mathbf{r}_{i})}\right]}_{\hat{H}_{0}}\\ &+\frac{1}{2}\frac{4\pi\hbar^{2}a_{sc}}{M}\left(\frac{1}{\sqrt{2\pi}\sigma}\right)^{3}\\ &\times\sum_{i\neq j}^{N}\exp^{-(\frac{1}{2\sigma^{2}})\{(r_{\perp i}-r_{\perp j})^{2}+(z_{i}-z_{j})^{2}\}}\end{split}$ (1)
where $\hat{H}_{0}=\sum_{i=1}^{N}\hat{h}(\mathbf{r}_{i})$ is the non-interacting part of the many-body Hamiltonian in the laboratory frame and
$\hat{L}_{z}^{\mathrm{lab}}=\sum_{i=1}^{N}\hat{l}_{z_{i}}^{\mathrm{lab}}=\frac{\hbar}{i}\sum_{i=1}^{N}(r_{i}\times\nabla_{i})_{z},$
$\hat{L}_{z}^{\mathrm{lab}}$ is the total angular momentum of the system in the laboratory frame. Here $r_{\perp}=\sqrt{x^{2}+y^{2}}$ is the radial distance of the particle from the trap axis; $\omega_{\perp}$ and $\omega_{z}$ are the confining harmonic frequencies in the $x$-$y$ plane and the axial direction, respectively. The first two terms of the many-body Hamiltonian in Eq. (1) comprises $\hat{H}_{0}$ and represent the kinetic energy and the external confining potential, while the last term describes the two-body interaction, modelled by a potential Gaussian in particle-particle separation. We introduce the anisotropic parameter $\lambda_{z}\equiv\frac{\omega_{z}}{\omega_{\perp}}\gg$1 which signifies that the trap is highly oblate spheroidal and hence the confined system is effectively quasi-2D. We choose $\hbar\omega_{\perp}$ as the unit of energy and define $a_{\perp}$$\sqrt{\frac{\hbar}{M\omega_{\perp}}}$ as the unit of length. In the Gaussian potential, $\sigma$ (scaled by $a_{\perp}$) is the interaction range of the potential. The effective two-body interaction strength in 2D, after tracing out the $z$-coordinate becomes
$g_{2}=\frac{4\pi\hbar^{2}a_{sc}}{M}\,\frac{1}{\sqrt{2\pi}}\sqrt{\frac{\lambda_{z}}{a_{\perp}^{2}\left[1+\left(\tfrac{\sigma}{a_{\perp}}\right)^{2}\lambda_{z}\right]}},$
which in the limit $\frac{\sigma}{a_{\perp}}\ll 1$ and in the dimensionless form reduces to
$g_{2}=\frac{4\pi a_{sc}}{a_{\perp}}\sqrt{\frac{\lambda_{z}}{2\pi}}$ (2)
where $a_{sc}$ is the s-wave scattering length for the two-body interaction. Experimentally, one can vary the scattering length $a_{sc}$ employing Feshbach resonance [43, 52]. In this study we assume $a_{sc}>0$ so that the effective interparticle potential is in the repulsive regime. When the interaction range approaches zero i.e $\sigma\to 0$, the Gaussian interaction reduces to the $\delta$-function potential $V(r_{\perp i},r_{\perp j})=g_{2}\delta(r_{\perp i}-r_{\perp j})$ [44].
The eigenspectrum for the many-body Hamiltonian $\hat{H}^{\mathrm{lab}}$ in Eq. (1), in given subspaces of total angular momentum $\hat{L}_{z}^{\mathrm{lab}}$, in the laboratory frame are obtained variationally by setting up the $N$-body eigenvalue equation
$\hat{H}^{\mathrm{lab}}\Psi(\mathbf{r}_{1},\mathbf{r}_{2},\ldots,\mathbf{r}_{N})=E\Psi(\mathbf{r}_{1},\mathbf{r}_{2},\ldots,\mathbf{r}_{N}),$ (3)
for the Hamiltonian in Eq. (1) with
$\Psi(\mathbf{r}_{1},\mathbf{r}_{2},\ldots,\mathbf{r}_{N})=\sum_{\nu}C_{\nu}\phi_{\nu}(\mathbf{r}_{1},\mathbf{r}_{2},\ldots,\mathbf{r}_{N}),$
where the set of parameters $\{C_{\nu}\}$ are to be determined variationally and the set $\{\phi_{\nu}(\mathbf{r}_{1},\mathbf{r}_{2},\ldots,\mathbf{r}_{N})\}$ are the $N$-body basis functions, satisfying the eigenvalue equation $\hat{H}_{0}\phi_{\nu}=E_{\nu}\phi_{\nu}$, where $H_{0}$ is the non-interacting part of the Hamiltonian in Eq. (1) and $\{\nu\}$ is the set of $N$-body quantum numbers. The $\phi_{\nu}$ is constructed as the symmetrized product of the harmonic oscillator single-particle basis functions in 2D satisfying the eigenvalue equation $\hat{h}(\mathbf{r})u_{n,m}(\mathbf{r})=\epsilon_{n,m}u_{n,m}(\mathbf{r})$ with
$\epsilon_{n,m}$ $=\left(\underbrace{2n_{r}+|m|}_{n}+1\right)\hbar\omega_{\perp}\equiv(n+1)\hbar\omega_{\perp},$
$\quad n=0,1,2,\ldots,\quad m=\underbrace{-n,-n+2,\ldots,n-2,+n}_{\text{in steps of 2}}.$
@@
$u_{n,m}(r_{\perp},\phi)=$ $\sqrt{\frac{\alpha_{\perp}^{2}}{\pi}\frac{\left(\tfrac{1}{2}[n-|m|]\right)!}{\left(\tfrac{1}{2}[n+|m|]\right)!}}\,e^{-\frac{\alpha_{\perp}^{2}r_{\perp}^{2}}{2}}\,e^{im\phi}(\alpha_{\perp}r_{\perp})^{|m|}$
$\times L_{\tfrac{1}{2}(n-|m|)}^{|m|}(\alpha_{\perp}^{2}r_{\perp}^{2}),\quad\alpha_{\perp}=\sqrt{\frac{M\omega_{\perp}}{\hbar}},$
$\hat{l}_{z}\,u_{n,m}(r_{\perp},\phi)=m\hbar\,u_{n,m}(r_{\perp},\phi).$
On carrying out the variational scheme, the $N$-body eigenvalue problem in Eq. (3) reduces to the matrix eigenvalue problem $\mathbf{H}\mathbf{C}=E\mathbf{C}$, where $\mathbf{C}$ is the column vector comprising of the variational parameters $\{C_{\nu}\}$ and $\mathbf{H}$ is the matrix representation of the Hamiltonian $\hat{H}^{\mathrm{lab}}$ in Eq. (1). The Hamiltonian matrix $\mathbf{H}$ is large but sparse and the lowest $100$ eigenvalues and eigenvectors are determined through iterative diagonalization [53]. The detailed scheme for finding the eigenspectrum can be found in Ref. [45].
Having obtained the eigenspectrum for the Hamitonian in Eq. (1) in given subspaces of the total angular momentum, we proceed to carry out the statistical analysis of the energy levels as described in the following.
@@ III.1 Nearest-neighbor spacing distribution
where $b$ is the Brody parameter and $a=[\Gamma(\frac{2+b}{1+b})]^{1+b}$ $\Gamma$ is Euler’s gamma function]. The parameter $b$ is like an interpolation parameter which for $b=0$ corresponds to the Poisson distribution while $b=1$ corresponds to the GOE distribution.
@@ III.2 The Dyson-Mehta $\Delta_{3}(L)$ statistic
To study the long-range correlations in the system, the Dyson-Mehta $\Delta_{3}(L)$ statistic is employed [5]. For the chosen energy interval $[a,a+L]$ of length $L$, the $\Delta_{3}$ statistic is defined by
$\Delta_{3}(a;L)=\frac{1}{L}\underset{A,B}{\min}\int_{a}^{a+L}[N(\epsilon)-A\epsilon-B]^{2}d\epsilon.$ (5)
@@
$\Delta_{3}(L)=\frac{L}{15}$
@@
$\Delta_{3}(L)=\frac{1}{\pi^{2}}\ln(L)-0.00695.$